# A Survey on 3D Gaussian Splatting

### v8

@@ 1. Introduction
The objective of image based 3D scene reconstruction is to convert a collection of views or videos capturing a scene into a digital 3D model that can be computationally processed, analyzed, and manipulated. This hard and long-standing problem is fundamental for machines to comprehend the complexity of real-world environments, facilitating a wide array of applications such as 3D modeling and animation, robot navigation, historical preservation, augmented/virtual reality, and autonomous driving.
The journey of 3D scene reconstruction began long before the surge of deep learning, with early endeavors focusing on light fields and basic scene reconstruction methods (Gortler et al., 2023; Levoy and Hanrahan, 2023; Buehler et al., 2023). These early attempts, however, were limited by their reliance on dense sampling and structured capture, leading to significant challenges in handling complex scenes and lighting conditions. The emergence of structure-from-motion (Snavely et al., 2006) and subsequent advancements in multi-view stereo (Goesele et al., 2007) algorithms provided a more robust framework for 3D scene reconstruction. Despite these advancements, such methods struggled with novel-view synthesis and texture loss. NeRF represents a quantum leap in this progression. By leveraging deep neural networks, NeRF enabled the direct mapping of spatial coordinates to color and density. The success of NeRF hinged on its ability to create continuous, volumetric scene functions, producing results with unprecedented fidelity. However, as with any burgeoning technology, this implementation came at a cost: i) Computational Intensity. NeRF based methods are computationally intensive (Garbin et al., 2021; Müller et al., 2022), often requiring extensive training times and substantial resources for rendering, especially for high-resolution outputs. ii) Editability. Manipulating scenes represented implicitly is challenging, since direct modifications to the neural network’s weights are not intuitively related to changes in geometric or appearance properties of the scene.
@@
The introduction of 3D GS is not just a technical advancement; it represents a fundamental shift in how we approach scene representation and rendering in computer vision and graphics. By enabling real-time rendering capabilities without compromising on visual quality, 3D GS opens up a plethora of possibilities for applications ranging from virtual reality
Figure 1. The number of published papers and official GitHub stars on 3D GS. The set of statistics is sourced from # Papers and # GitHub Stars.
and augmented reality to real-time cinematic rendering and beyond (Kalkofen et al., 2008; Patney et al., 2016; Albert et al., 2017; Jiang et al., 2024c). This technology holds the promise of not only enhancing existing applications but also enabling new ones that were previously unfeasible due to computational constraints. Furthermore, 3D GS’s explicit scene representation offers unprecedented flexibility to control the objects and scene dynamics, a crucial factor in complex scenarios involving intricate geometries and varying lighting conditions (Lu et al., 2024b; Saito et al., 2024; Zhang et al., 2024e). This level of editability, combined with the efficiency of the training and rendering process, positions 3D GS as a transformative force in shaping future developments in relevant fields.
In an effort to assist readers in keeping pace with the swift evolution of 3D GS, we provide the first survey on 3D GS, which presents a systematic and timely collection of the most significant literature on the topic. Given that 3D GS is a very recent innovation (c.f. Fig. 1), this survey focuses in particular on its principles, and the diverse developments and contributions that have emerged since its introduction. We systematically review the foundational period of 3D GS, with selected works primarily sourced from top-tier conferences. Our analysis concentrates on the initial explosion of research following its introduction in 2023 through 2024, covering the theoretical underpinnings, landmark developments, and early applications that shaped the field. Acknowledging the nascent yet rapidly evolving nature of 3D GS, this survey is inevitably a biased view, but we strive to offer a balanced perspective that reflects both the current state and the future potential of this field. Our aim is to encapsulate the primary research trends and serve as a valuable resource for both researchers and practitioners eager to understand and contribute to this rapidly evolving domain. The distinctions of this survey from existing literature (Fei et al., 2024; Dalal et al., 2024; Bao et al., 2024; Wu et al., 2024e) are evident in the following aspects:
@@ 2. Background
In this section, we first provide a brief formulation of radiance fields (Sec. 2.1), including both implicit and explicit ones. Sec. 2.2 further establishes linkages with relevant rendering algorithms and terminologies. For a comprehensive overview of radiance fields, scene reconstruction and representation, and rendering methods, please see the excellent surveys (Kobbelt and Botsch, 2004; Xie et al., 2022; Wang et al., 2024b; Tewari et al., 2022; Han et al., 2019) for more insights.
§1§2§3§4§5§6§7§8§2.1§2.2§3.1§3.2§4.1§4.2§4.3§4.4§4.5§4.6§4.7§5.1§5.2§5.3§5.4§5.5§5.6§5.7§6.1§6.2§6.3§6.4§6.5
@@ 2.1. Radiance Field
$\bullet$ Implicit Radiance Field. An implicit radiance field represents light distribution in a scene without explicitly defining the geometry of the scene. In the deep learning era, neural networks are often used to learn a continuous volumetric scene representation (Mescheder et al., 2019; Park et al., 2019). The most prominent example is NeRF (Mildenhall et al., 2020). In NeRF (Fig. 4a), one or more MLPs are used to map a set of spatial coordinates $(x,y,z)$ and viewing directions $(\theta,\phi)$ to color $c$ and volume density $\sigma$:
(1) $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(c,\sigma)\leftarrow\text{MLP}(x,y,z,\theta,\phi)$}.
This format allows for a differentiable and compact representation of complex scenes, albeit often at the cost of high computational load due to volumetric ray marching. Note that typically, the color $c$ is direction-dependent, whereas the volume density $\sigma$ is not (Mildenhall et al., 2020).
$\bullet$ Explicit Radiance Field. An explicit radiance field directly represents the distribution of light in a discrete spatial structure, such as a voxel grid or a set of points (Sun et al., 2022; Fridovich-Keil et al., 2023). Each element in this structure stores the radiance information for its respective location. This allows for direct and often faster access to radiance data but at the cost of higher memory usage and potentially lower resolution. Similar to the implicit radiance field, the explicit one is written as:
(2) $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(c,\sigma)\leftarrow\text{DataStructure}(x,y,z,\theta,\phi)$},
@@
$\bullet$ 3D Gaussian Splatting: Best-of-Both Worlds. 3D GS (Kerbl et al., 2023) is an explicit radiance field with the advantages of implicit radiance fields. Concretely, it leverages the strengths of both paradigms by utilizing learnable 3D Gaussians as the basis elements of DataStructure. Note that 3D GS encodes the opacity $\alpha$ directly for each Gaussian, as opposed to approaches of first establishing density $\sigma$ and then computing opacity based on that density. As in previous reconstruction work, 3D Gaussians are optimized under the supervision of multi-view images to represent the scene. Such a 3D Gaussian based differentiable pipeline combines the benefits of neural network based optimization and explicit, structured data storage. This hybrid approach aims to achieve real-time, high-quality rendering and requires less training time, particularly for complex scenes and high-resolution outputs.
@@ 2.2. Context and Terminology
(3) $\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}C(\bm{r})=\int_{t_{\text{near}}}^{t_{\text{far}}}T(t)\,\sigma(\bm{r}(t))\,c(\bm{r}(t),\bm{d})\,dt,$}
@@
$\bullet$ Point-based rendering represents another class of rendering algorithms, of which 3D GS introduces a notable implementation. Its simplest form (Grossman and Dally, 1998) rasterizes point clouds with a fixed size, which introduces drawbacks such as holes and rendering artifacts. Seminal works addressed these limitations through various methods, including: i) splatting point primitives with a spatial extent (Zwicker et al., 2001b, a, 2002; Ren et al., 2002), and ii) more recently, embedding neural features directly into points for subsequent network-based rendering (Aliev et al., 2020; Rückert et al., 2022). 3D GS uses 3D Gaussian as the point primitive that contains explicit attributes (e.g., color and opacity) instead of implicit neural features. The rendering approach, i.e., point-based $\alpha$-blending (exemplified in Eq. 5), shares the same image formation model as NeRF-style volumetric rendering (Eq. 3) (Kerbl et al., 2023), but demonstrates substantial speed advantages. This advantage originates from fundamental algorithmic differences. NeRFs approximate a line integral along a ray for each pixel, requiring expensive sampling. Point-based methods render point clouds using rasterization, which inherently benefits from parallel computational strategies (Lassner and Zollhofer, 2021).
@@
3.1. Rendering with Learned 3D Gaussians
Consider a scene represented by (millions of) optimized 3D Gaussians. The objective is to generate an image from a specified camera pose. Recall that NeRFs approach this task through computationally demanding volumetric ray-marching,
$x,y,z,\theta,\phi$
$c\penalty 10000\ \&\penalty 10000\ \sigma$
Figure 4. NeRFs vs. 3D GS. (a) NeRF samples along the ray and then queries the MLP to obtain colors and densities, which can be seen as a backward mapping (ray tracing). (b) 3D GS projects all 3D Gaussians into the image space (i.e., splatting) and then performs parallel rendering, which can be viewed as a forward mapping (rasterization).
sampling 3D space points per pixel. Such a paradigm struggles with high-resolution image synthesis, failing to achieve real-time rendering, especially for platforms with limited computing resources (Kerbl et al., 2023). By contrast, 3D GS begins by projecting these 3D Gaussians onto a pixel-based image plane, a process termed “splatting” (Zwicker et al., 2001a, 2002) (c.f. Fig. 4b). Afterwards, 3D GS sorts these Gaussians and computes the value for each pixel. As shown in Fig. 4, the rendering of NeRFs and 3D GS can be viewed as an inverse process of each other. In what follows, we begin with the definition of a 3D Gaussian, which is the minimal element of the scene representation in 3D GS. Next, we describe how these 3D Gaussians can be used for differentiable rendering. Finally, we introduce the acceleration technique used in 3D GS, which is the key to fast rendering.
$\bullet$ Properties of 3D Gaussian. A 3D Gaussian is characterized by its center (position) $\bm{\mu}$, opacity $\alpha$, 3D covariance matrix $\bm{\Sigma}$, and color $c$. $c$ is represented by spherical harmonics for view-dependent appearance. All the properties are learnable and optimized through back-propagation.
@@ 3.1. Rendering with Learned 3D Gaussians
$\bullet$ Splatting (Fig. 3a). In this step, 3D Gaussians (ellipsoids) in 3D space are projected into 2D image space (ellipses). The projection proceeds through two transformations: first, transforming 3D Gaussians from world coordinates to camera coordinates using the viewing transformation, and subsequently splatting these Gaussians into 2D image space via an approximation of the projective transformation. Mathematically, given the 3D covariance matrix $\bm{\Sigma}$ describing a 3D Gaussian’s spatial distribution, and the viewing transformation matrix $\bm{W}$, the 2D covariance matrix $\bm{\Sigma}^{\prime}$ characterizing the projected 2D Gaussian is computed through:
@@
$\bullet$ Rendering by Pixels. Before delving into the final version of 3D GS which utilizes several techniques to boost parallel computation, we first elaborate on its simpler form to offer insights into its basic working mechanism. Given the position of a pixel $\bm{x}$, its distance to all overlapping Gaussians, i.e., the depths of these Gaussians, can be computed through the viewing transformation matrix $\bm{W}$, forming a sorted list of Gaussians $\mathcal{N}$. Then, $\alpha$-blending is adopted to compute the final color of this pixel:
(5) $C=\sum_{n=1}^{|\mathcal{N}|}c_{n}\alpha^{\prime}_{n}\prod_{j=1}^{n-1}\left(1-\alpha^{\prime}_{j}\right),$
where $c_{n}$ is the learned color. The final opacity $\alpha^{\prime}_{n}$ is the multiplication result of the learned opacity $\alpha_{n}$ and the Gaussian, defined as follows:
(6) $\alpha^{\prime}_{n}=\alpha_{n}\times\text{exp}\big(-\frac{1}{2}(\bm{x}^{\prime}-\bm{\mu}^{\prime}_{n})^{\top}\bm{\Sigma}^{\prime-1}_{n}(\bm{x}^{\prime}-\bm{\mu}^{\prime}_{n})\big),$
@@
5
Figure 5. An illustration of the tile based parallel (at the pixel-level) rendering. All the pixels within a tile (Tile1 here) access the same ordered Gaussian list stored in a shared memory for rendering. As the system processes each Gaussian sequentially, every pixel in the tile evaluates the Gaussian’s contribution according to the distance (i.e., the $\exp$ term in Eq. 6). Therefore, the rendering for a tile can be completed by iterating through the list of Gaussians just once. The computation for the red Gaussian follows a similar way and is omitted here for simplicity.
$\bullet$ Parallel Rendering. After replication, 3D GS combines the respective tile ID with the depth value obtained from the view transformation for each Gaussian. This results in an unsorted list of bytes where the upper bits represent the tile ID and the lower bits signify depth. By doing so, the sorted list can be directly utilized for rendering (i.e., alpha compositing). Fig. 3c and Fig. 3d provide the visual demonstration of such concepts. It’s worth highlighting that rendering each tile and pixel occurs independently, making this process highly suitable for parallel computations. An additional benefit is that each tile’s pixels can access a common shared memory and maintain an uniform read sequence (Fig. 5), enabling parallel execution of alpha compositing with increased efficiency. In the official implementation of the original paper (Kerbl et al., 2023), the framework regards the processing of tiles and pixels as analogous to the blocks and threads, respectively, in CUDA programming architecture.
In a nutshell, 3D GS introduces several approximations during rendering to enhance computational efficiency while maintaining a high standard of image synthesis quality.

### v9

@@ 1. Introduction
The objective of image based 3D scene reconstruction is to convert a collection of views capturing a scene into a digital 3D model that can be computationally processed, analyzed, and manipulated. This hard and long-standing problem is fundamental for machines to comprehend real-world environments, facilitating a wide array of applications such as 3D modeling and animation, robot navigation, historical preservation, augmented/virtual reality, and autonomous driving.
The journey of 3D scene reconstruction began long before the surge of deep learning, with early endeavors focusing on light fields and basic reconstruction methods (Gortler et al., 2023; Levoy and Hanrahan, 2023; Buehler et al., 2023). These early attempts, however, were limited by their reliance on dense sampling and structured capture, leading to challenges in handling complex scenes and lighting conditions. The emergence of structure-from-motion (Snavely et al., 2006) and subsequent advancements in multi-view stereo (Goesele et al., 2007) algorithms provided a more robust framework. Despite these advancements, such methods struggled with novel-view synthesis and texture loss. NeRF represents a quantum leap in this progression. By leveraging neural networks, NeRF enabled the direct mapping of spatial coordinates to color and density. The success of NeRF hinged on its ability to create continuous, volumetric scene functions, producing results with unprecedented fidelity. However, as with any burgeoning technology, it came at a cost: i) Computational Intensity. NeRF based methods are computationally intensive (Garbin et al., 2021; Müller et al., 2022), often requiring substantial training times and resources for rendering, especially for high-resolution outputs. ii) Editability. Manipulating scenes represented implicitly is challenging, since direct modifications to the neural network’s weights are not intuitively related to changes in geometric or appearance properties of the scene.
@@
The introduction of 3D GS is not just a technical advancement; it represents a fundamental shift in how we approach scene representation and rendering in computer vision and graphics. By enabling real-time rendering capabilities without compromising on visual quality, 3D GS opens up a plethora of possibilities for applications ranging from virtual reality and augmented reality to real-time cinematic rendering and beyond (Kalkofen et al., 2008; Patney et al., 2016; Albert et al., 2017; Jiang et al., 2024a). This technology holds the promise of not only enhancing existing applications but also enabling new ones that were previously unfeasible due to computational constraints. Furthermore, 3D GS’s explicit scene representation offers unprecedented flexibility to control the objects and scene dynamics, a crucial factor in complex scenarios involving intricate geometries and varying lighting conditions (Lu et al., 2024b; Saito et al., 2024; Zhang et al., 2024i). This level of editability, combined with the efficiency of the training and rendering process, positions 3D GS as a transformative force in shaping future developments in relevant fields.
In an effort to assist readers in keeping pace with the swift evolution of 3D GS, we provide the first survey on 3D GS, which presents a systematic and timely collection of the most significant literature on the topic. Given that GS is a very recent innovation (c.f. Fig. S1), this survey focuses in particular on its principles, and the diverse developments and contributions that have emerged since its introduction. We systematically review the foundational period of GS, with selected works primarily sourced from top-tier conferences. Our analysis concentrates on the initial explosion of research following its introduction in 2023 through 2024, covering the theoretical underpinnings, landmark developments, and early applications that shaped the field. Acknowledging the nascent yet rapidly evolving nature, this survey is inevitably a biased view, but we strive to offer a balanced perspective that reflects both the current state and the future potential of this field. Our aim is to encapsulate the primary research trends and serve as a valuable resource for both researchers and practitioners eager to understand and contribute to this rapidly evolving domain. The distinctions of this paper from existing surveys (general (Fei et al., 2024; Dalal et al., 2024; Bao et al., 2024; Wu et al., 2024f) or domain-specific (Bagdasarian et al., 2025; Ali et al., 2025; He et al., 2025)) are evident in the following aspects:
@@ 2. Background
This section provides a brief formulation of radiance fields (Sec. 2.1), including both implicit and explicit ones. Sec. 2.2 further establishes linkages with relevant rendering algorithms and terminologies. Please see (Kobbelt and Botsch, 2004; Xie et al., 2022; Wang et al., 2024d; Tewari et al., 2022; Han et al., 2019) for a comprehensive overview of radiance fields, scene reconstruction and representation, and rendering methods.
§1§2§3§4§5§6§7§8§2.1§2.2§3.1§3.2§4.1§4.2§4.3§4.4§4.5§4.6§4.7§4.8§5.1§5.2§5.3§5.4§5.5§5.6§5.7§6.1§6.2§6.3§6.4§6.5
@@ 2.1. Radiance Field
$\bullet$ Implicit Radiance Field. An implicit radiance field represents light distribution in a scene without explicitly defining the geometry of the scene. In the deep learning era, neural networks are often used to learn a continuous volumetric scene representation (Mescheder et al., 2019; Park et al., 2019). The most prominent example is NeRF (Mildenhall et al., 2020). In NeRF (Fig. S2a), one or more MLPs are used to map a set of spatial coordinates $(x,y,z)$ and viewing directions $(\theta,\phi)$ to color $c$ and volume density $\sigma$:
(1) $(c,\sigma)\leftarrow\text{MLP}(x,y,z,\theta,\phi).$
This format allows for a differentiable and compact representation of scenes, albeit often at the cost of high computational load. Note that typically, the color $c$ is direction-dependent, whereas the volume density $\sigma$ is not (Mildenhall et al., 2020).
$\bullet$ Explicit Radiance Field. An explicit radiance field directly represents the distribution of light in a discrete spatial structure, such as a voxel grid or a set of points (Sun et al., 2022; Fridovich-Keil et al., 2023). Each element in this structure stores the radiance information for its respective location. This allows for direct and often faster access to radiance data but at the cost of higher memory usage and potentially lower resolution. Plenoxels (Fridovich-Keil et al., 2022) highlighted this trend by discarding the MLP and directly optimizing a sparse voxel grid with spherical harmonic coefficients, showing that a fully explicit grid can achieve NeRF-level quality while converging much faster. Similar to the implicit radiance field, the explicit one is written as:
(2) $(c,\sigma)\leftarrow\text{DataStructure}(x,y,z,\theta,\phi),$
@@
$\bullet$ 3D Gaussian Splatting: Best-of-Both Worlds. 3D GS (Kerbl et al., 2023) is an explicit radiance field with the advantages of implicit radiance fields. Concretely, it uses the strengths of both paradigms by utilizing learnable 3D Gaussians as the basis elements of DataStructure. Note that 3D GS encodes the opacity $\alpha$ directly for each Gaussian, as opposed to approaches of first establishing density $\sigma$ and then computing opacity based on that density. As in previous reconstruction work, 3D Gaussians are optimized under the supervision of multi-view images to represent the scene. Such a 3D Gaussian based differentiable pipeline combines the benefits of neural network based optimization and explicit, structured data storage. This hybrid approach aims to achieve real-time, high-quality rendering and requires less training time, particularly for complex scenes and high-resolution outputs. In addition, from a differentiable rendering perspective, 3D GS complements mesh-based differentiable rasterizers such as nvdiffrast (Laine et al., 2020), which operate on explicit triangle meshes using high-performance modular rasterization primitives.
@@ 2.2. Context and Terminology
(3) $C(\bm{r})=\int_{t_{\text{near}}}^{t_{\text{far}}}T(t)\,\sigma(\bm{r}(t))\,c(\bm{r}(t),\bm{d})\,dt,$
@@
$\bullet$ Point-based rendering represents another class of rendering algorithms, of which 3D GS introduces a notable implementation. Its simplest form (Grossman and Dally, 1998) rasterizes point clouds with a fixed size, which introduces drawbacks such as holes and rendering artifacts. Seminal works addressed these limitations through various methods, including: i) employing multi-resolution level-of-detail hierarchies like QSplat (Rusinkiewicz and Levoy, 2000) to handle large-scale scenes, ii) splatting point primitives with a spatial extent (Zwicker et al., 2001b, a, 2002; Ren et al., 2002), and iii) more recently, embedding neural features directly into points for subsequent network-based rendering (Aliev et al., 2020; Rückert et al., 2022). 3D GS uses 3D Gaussian as the point primitive that contains explicit attributes (e.g., color and opacity) instead of implicit neural features. The rendering approach, i.e., point-based $\alpha$-blending (exemplified in Eq. 5), shares the same image formation model as NeRF-style volumetric rendering (Eq. 3) (Kerbl et al., 2023), but demonstrates substantial speed advantages. This advantage originates from fundamental algorithmic differences. NeRFs approximate a line integral along a ray for each pixel, requiring expensive sampling. Point-based methods render point clouds using rasterization, which inherently benefits from parallel computational strategies (Lassner and Zollhofer, 2021).
@@
3.1. Rendering with Learned 3D Gaussians
Consider a scene represented by optimized 3D Gaussians. The objective is to generate an image from a specified camera pose. Recall that NeRFs approach this task through volumetric ray-marching, sampling 3D space points per pixel. Such a paradigm struggles with high-resolution image synthesis, failing to achieve real-time rendering, especially for platforms with limited computing resources (Kerbl et al., 2023). By contrast, 3D GS begins by projecting these 3D Gaussians onto a pixel-based image plane, a process termed “splatting” (Zwicker et al., 2001a, 2002) (c.f. Fig. S2b). Afterwards, 3D GS sorts these Gaussians and computes the value for each pixel. As shown in Fig. S2, the rendering of NeRFs and 3D GS can be viewed as an inverse process of each other. In what follows, we begin with the definition of a 3D Gaussian, which is the minimal element of the scene representation in 3D GS. Next, we describe how these 3D Gaussians can be used for differentiable rendering. Finally, we introduce the acceleration technique used in 3D GS, which is the key to fast rendering.
@@ 3.1. Rendering with Learned 3D Gaussians
$\bullet$ Properties of 3D Gaussian. A 3D Gaussian is characterized by its center $\bm{\mu}$, opacity $\alpha$, 3D covariance matrix $\bm{\Sigma}$, and color $c$. $c$ is represented by spherical harmonics (SH) coefficients to model view-dependent appearance. Unlike a static RGB value, SH functions allow the color to vary based on the viewing angle. This enables the representation to capture non-Lambertian effects, such as specular highlights and lighting variations, by evaluating the SH coefficients with the specific view direction during rendering. All the properties are learnable and optimized through back-propagation.
@@
$\bullet$ Splatting (Fig. 2a). In this step, 3D Gaussians (ellipsoids) in 3D space are projected into 2D image space (ellipses). The projection proceeds through two transformations: first, transforming 3D Gaussians from world coordinates to camera coordinates using viewing transformation, and subsequently splatting these Gaussians into 2D image space via an approximation of projective transformation. Given the 3D covariance matrix $\bm{\Sigma}$ describing a 3D Gaussian and viewing transformation matrix $\bm{W}$, the 2D covariance matrix $\bm{\Sigma}^{\prime}$ characterizing the projected 2D Gaussian is computed through:
@@
$\bullet$ Rendering by Pixels. Before delving into the final version of GS with highly parallel computation, we first elaborate on its simpler form to offer insights into its basic working mechanism. Given the position of a pixel $\bm{x}$, its distance to all overlapping Gaussians, i.e., the depths of these Gaussians, can be computed through the viewing transformation matrix $\bm{W}$, forming a sorted list of Gaussians $\mathcal{N}$. Then, $\alpha$-blending (Porter and Duff, 1984) is adopted to compute the final color of this pixel:
(5) $C=\sum\nolimits_{n=1}^{|\mathcal{N}|}c_{n}\alpha^{\prime}_{n}\prod\nolimits_{j=1}^{n-1}\left(1-\alpha^{\prime}_{j}\right),$
where $c_{n}$ is the learned color. The final opacity $\alpha^{\prime}_{n}$ is the multiplication result of the learned opacity $\alpha_{n}$ and the Gaussian:
(6) $\alpha^{\prime}_{n}=\alpha_{n}\times\text{exp}\big(-0.5(\bm{x}^{\prime}-\bm{\mu}^{\prime}_{n})^{\top}\bm{\Sigma}^{\prime-1}_{n}(\bm{x}^{\prime}-\bm{\mu}^{\prime}_{n})\big),$
@@
$\bullet$ Parallel Rendering. After replication, 3D GS combines the respective tile ID with the depth value obtained from the view transformation for each Gaussian. This results in an unsorted list of bytes where the upper bits represent the tile ID and the lower bits signify depth. By doing so, the sorted list can be directly utilized for rendering (i.e., alpha compositing). Fig. 2c and Fig. 2d provide the visual demonstration of such concepts. It’s worth highlighting that rendering each tile and pixel occurs independently, making this process highly suitable for parallel computations. An additional benefit is that each tile’s pixels can access a common shared memory and maintain an uniform read sequence (Fig. S3), enabling parallel execution of alpha compositing with increased efficiency. In the official implementation of the original paper (Kerbl et al., 2023), the framework regards the processing of tiles and pixels as analogous to the blocks and threads, respectively, in CUDA programming architecture.